OpenStudy (anonymous):

Write each fraction as a sum or difference Please also explain how to do it, I'm very confused 2x+3 ---- 3

OpenStudy (anonymous):

is it \[\frac{ 2x+3 }{ 3}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

now divide 3 to both 2x and 3 you will get this, \[\frac{ 2x }{3 } + \frac{ 3 }{ 3 } = \frac{ 2x }{ 3 } + 1\]

OpenStudy (anonymous):

Thanks, also. Whats \[\frac{5t-12 }{ 10 }\]